Three different companies each purchased trucks on January 1, 2018, for $54,000. Each truck was expected to last four years or 250,000 miles. Salvage value was estimated to be $5,000. All three trucks were driven 76,000 miles in 2018, 65,000 miles in 2019, 42,000 miles in 2020, and 71,000 miles in 2021. Each of the three companies earned $43,000 of cash revenue during each of the four years. Company A uses straight-line depreciation, company B uses double-declining-balance depreciation, and company C uses units-of-production depreciation.

Answer each of the following questions. Ignore the effects of income taxes.

b-1. Calculate the net income for 2021? (Round "Per Unit Cost" to 3 decimal places.)

Company A:

Company B:

Company C:

b-2. Which company will report the lowest amount of net income for 2021?

c-1. Calculate the book value on the December 31, 2020, balance sheet? (Round "Per Unit Cost" to 3 decimal places.)

Company A:

Company B:

Company C:

c-2. Which company will report the highest book value on the December 31, 2020, balance sheet?

d-1. Calculate the retained earnings on the December 31, 2021, balance sheet?

Company A:

Company B:

Company C:

d-2. Which company will report the highest amount of retained earnings on the December 31, 2021, balance sheet?
Answer #1

b-1:

Net Income
Company A $ 30,750
Company B 41,250
Company C 29,868

Company A : Straight-line Depreciation Expense = $ ( 54,000 - 5,000 ) / 4 = $ 12,250. Net Income : $ 43,000 - $ 12,250 = $ 30,750

Company B: Double-declining Depreciation Expense = $ 54,000 * 50 % * 50 % * 50 % * - 5,000 = $ 1,750.

Double declining rate of depreciation = 100 / 4 * 2 = 50 %.

Total accumulated depreciation over the useful life of the asset cannot exceed $ 49,000 in this case.

Net Income : $ 43,000 - $ 1,750 = $ 41,250.

Company C : UOP Depreciation Expense = $ 49,000 / 250,000 miles * ( 67,000) = $ 13,132.

Net Income = $ 43,000 - $ 13,132 = $ 29,868.

Total miles driven cannot exceed the initially estimated 250,000 miles. ( 76,000 + 65,000 + 42,000 + 67,000)

b-2. : Company C will report the lowest amount of net income.

c-1:

Book Value
Company A $ 17,250
Company B 6,750
Company C 18,132

Company A: Book value = $ 54,000 - ( $ 12,250 x 3 ) = $ 17,250.

Company B: Book value : $ 54,000 x 50 % x 50 % x 50 % = $ 6,750

Company C: Book value : $ 54,000 - ( 76,000 + 65,000 + 42,000) * $ 49,000 / 250,000 = $ 18,132.

c-2: Company C will report the highest book value on the December 31, 2020 balance sheet.

d-1:

Retained Earnings, December 31, 2021
Company A $ 123,000
Company B 123,000
Company C 123,000

Total cash revenue earned over 4 years = $ 43,000 x 4 = $ 172,000.

Retained earnings balance as of December 31, 2021 = Total cash earnings- Accumulated Depreciation = $ 172,000 - $ 49,000 = $ 123,000.

d-2.: All three companies will report the same retained earnings on the December 31, 2021 balance sheet.